The area under the normal curve between Z = 1 and Z = 2 is ________________ the area under the normal curve between Z = 0 and Z = 1. A. Less than B. Greater than C. Equal to D. A, B or C above dependent on the value of the mean E. A, B or C above dependent on the value of the standard deviation
Answer #1

P(1 < Z < 2)

= P(Z < 2) - P(Z < 1)

= 0.9772 - 0.8413

= 0.1359

P(0 < Z < 1)

= P(Z < 1) - P(Z < 0)

= 0.8413 - 0.5

= 0.3413